# Data Quality Issues in Flare Prediction using Machine Learning Models

### v1

@@ 4 Comparison of Data Products from a Flare Prediction Perspective
In this section, we compare the SWPC-FTP operational flare list with the NCEI Science-Quality list by evaluating the numerical skill scores of machine learning models trained on definitive or near-real-time SHARP summary vector parameters. Note that we do not exhaust the available machine learning models and do not aim to optimize reported skill scores here, since, as discussed in the Introduction, (i) when properly trained, commonly used machine learning models do not generate significantly different results, and (ii) the skill scores of machine learning models depend heavily on sample selection and sample split schemes. Therefore, we use common practice for the choice of machine learning model and sample split here to demonstrate the differences in the various data products’ effects on the trained model’s performance.
@@ 4.1 Data Preparation and Model Setting
To assess how data quality affects flare forecasting performance, we consider two sources of flare labels (the SWPC-FTP list and the Science-Quality list) and two versions of predictor data (HMI near-real-time SHARPs and definitive calibrated SHARPs), see Table 9. We evaluate two representative model families: a deep learning architecture (LSTM) and a conventional statistical model (logistic regression with PCA, i.e., principal component analysis). All models are trained on Solar Cycle 24 (2010-05-01 to 2019-12-31) and tested independently on years 2020–2024. The task is formulated as the prediction of $\geq$ M-class flares in the next 24, 12, 6 hours.
@@
We follow the data preparation and sample construction methods described in  (Jiao2020lstm). The input data consists of 24-hour time series extracted from the full time series of SHARP parameters of ARs. To ensure data quality, some sequences are excluded, particularly when the ARs are located near the limb. The criteria for excluding unqualified time sequences are as follows.
@@
In this paper, we focus on a binary classification problem that distinguishes strong(M/X) from Weak (A/B) flares. We exclude the C-class category because its physical significance is ambiguous. C-class flares may be comparable to either small M-class flares or strong B-class flares, making it an unreliable separator boundary and thus vague the training process (chen2019solar). In this study, we adopt two models: an LSTM model with a sigmoid output layer and a logistic regression model. The LSTM-based model is widely used in solar flare forecasting for handling time-series inputs ((chen2019iden), (Liu2019lstm), (Wang2020lstm)) and has been shown to outperform other machine learning models ((Wang2020lstm)). The network consists of a two-layer LSTM with a hidden size of 30 units and dropout of 0.3 between layers, followed by a ReLU activation and a fully connected layer that maps the final (or truncated) hidden representation to a single scalar logit. This logit value is interpreted as the pre-sigmoid score for the probability of a strong flare.
We also include logistic regression due to its simplicity, which in fact results in stability for trained models, and interpretability. Given the $k$-hour forecasting window used to extract SHARP parameters, the input of the LSTM model for each target is a $5k\times 20$ matrix, where each row represents a 20-dimensional vector of SHARP parameters. For the logistic model, we first flatten each input along the time dimension into a $100k\times 1$ vector.
@@
    For the logistic regression model, we then apply principal component analysis (PCA) to reduce dimensionality, retaining 98% variance. For the 24-hr forecasting window, 20 principal components already cover over 98% variance. The projection coefficients for these principal components are derived from the training set.
@@
    The LSTM model is optimized using the Adam optimizer with a binary cross-entropy loss.
@@
    The logistic regression model is fitted with $L_{2}$ regularization. The penalty strength is optimized with 5-fold cross-validation and the ROC-AUC scoring objective.
After each bootstrap model is trained, we evaluate probabilistic predictions on the held-out validation set and determine the optimal decision threshold that maximizes the True Skill Statistic (TSS) over a grid of candidate thresholds in $(0,1)$. The optimal threshold and corresponding validation TSS scores are recorded for every bootstrap sample. To properly evaluate the performance of the machine learning algorithms, we split the samples (time sequences of SHARP parameters and their corresponding maximum flare magnitudes) into training and testing sets. Note that splitting by Active Region (AR) or HARP is crucial to prevent information leakage from the training set to the testing set. In this study, all training/testing splits are performed chronologically by HARP, which mimics the operational setting but may not result in the highest skill scores. See wang2020solar for detailed discussions on sample splitting schemes and their impacts on flare prediction results.
It is important to note that varying definitions of positive and negative samples make direct comparisons of results from existing literature challenging, as discussed in chen2024solar. The purpose of this paper is not to propose a universal method for sample construction but rather to compare model performance across different datasets. We also demonstrate numerically that the differences arising from data inconsistency or quality are not affected by the choice of the sample construction method.
@@ 4.2 Science-Quality versus Operational “Response”
To quantify the effect of flare-label quality on the prediction models’ skill scores, we compare the LSTM model and the logistic regression model for 6/12/24 forecasting trained using Science-Quality and SWPC-FTP flare list while fixing the SHARPs parameter inputs at definitive SHARPs. Box plots of six commonly adopted skill scores (TSS, HSS, POD, F1, FAR, ACC, see, e.g., chen2019iden for definitions) for the 24-hour forecasting windows are shown in Figure 13 and Figure 14 for the LSTM model and the logistic regression model, respectively. The 6-hour and 12-hour window prediction results exhibit consistent trends as these. The numerical results for all different model setups are given in Tables in Appendix B.
@@
We trained models using data from Solar Cycle 24 (5/1/2010–12/31/2019) and tested them with different time intervals within Solar Cycle 25. The evaluation spans four distinct testing intervals that correspond to different phases of Solar Cycle 25. The 2020–2021 period reflects the early rising phase and is near solar minimum, characterized by extremely low flare occurrence and a highly imbalanced dataset. The 2022 period represents an evolving phase with increasing magnetic complexity and flare frequency. The 2023–2024 interval corresponds to the solar maximum, featuring intense flare activity and a high positive sample ratio (often exceeding 0.8). Finally, the combined 2020–2025 evaluation window averages performance across the full range of solar activity and reflects the overall performance of the model with a balanced test set.
@@
Across the test periods when the solar activity is relatively low, the models trained with SWPC-FTP operational flare list achieve higher predictive skill scores than those trained with Science-Quality labels. This improvement is clearer for logistic models. During the solar minimum (20200101-20211231), the logistic regression model improves from a TSS of 0.15 [0.11,0.18] (Science-Quality) to 0.37 [0.35,0.38] (SWPC-FTP), and the LSTM model improves from 0.42 [0.16,0.59] to 0.57 [0.40,0.69]. Similar margins in performance win appear in other metrics. However, for the LSTM model, the difference is not substantial. However, we do observe more stability with the logistic regression model results (with consistently shorter boxes in Figure 14 as compared to Figure 13) as compared to LSTM results. During the high-activity period (2023–2024) when the positive-sample ratio exceeds 0.85, the Science-Quality-trained models achieve better performance except for the TSS score of the logistic models. Across all the test periods, models trained with SWPC-FTP list get lower FAR scores, demonstrating that models trained with operational flare lists are better at identifying flares without increasing false alarms.
@@
Figure 13: Box plots of six skill scores for different testing year choices (given in the x-axis) in the solar cycle 25 of the M/X versus A/B classification task by the LSTM model trained on definitive SHARPs as inputs with a 24-hour forecasting window. The blue boxes represent the Science-Quality flare list, while the orange boxes represent the SWPC-FP operational flare list. The dashed lines indicate the positive sample proportions in the test set, corresponding to the Y axis on the right-hand-side of each sub-figure. Note that a smaller FAR means better performance. On each box, the black line marks the mean of the 30 independent bootstrap runs, and the triangle marks the median. The lower and upper bounds of the boxes correspond to the first and third quartiles $Q_{1}$ and $Q_{3}$. The upper and lower error bars are at $Q_{3}+1.5(Q_{3}-Q_{1})$ and $Q_{3}-1.5(Q_{3}-Q_{1})$, respectively. The small dots outside of the boxes show the data points outside the error bars (outliers). The mean value and median are calculated including the outliers.
@@
During solar minimum, LSTM models trained with the Science-Quality flare lists achieve higher TSS and F1 scores than the logistic regression models trained on the same labels. For the SWPC-FTP lists, the LSTM models perform slightly better or about the same as the logistic models. However, during solar maximum, when trained on identical flare lists, the LSTM models perform worse than the logistic models. In addition, the performance differences between LSTM and logistic models are noticeably larger during solar maximum than during solar minimum. This suggests that the more complex flare environment during solar maximum makes model choice more critical and that simpler models as logistic regression can have an advantage.
@@ 4.3 Near-real-time versus Definitive SHARP
We begin by examining the forecasting model performances when using the DEF-SHARPs and NRT-SHARPs datasets. Although the HARP numbers differ between the two products, we find that the conditional flare-response distributions (given the NOAA AR number) are highly similar. For each SHARP observation (every 12 minutes), we match flares from the Science-Quality list using the NOAA active-region number and flare peak time. Observations associated with M- or X-class flares are labeled as positive, and those associated with A- or B-class flares are labeled as negative, consistent with our model formulation. Figure 16 shows the yearly positive rate and positive-sample counts for both DEF-SHARPs and NRT-SHARPs. Despite having fewer valid observations, the NRT-SHARPs product exhibits a positive-rate pattern that closely matches the definitive product across all training years.
@@
Figure 17 presents the average TSS and F1 score gains across eight model configurations. Because the NRT-SHARPs dataset is extremely imbalanced during 2023–2024, we compare performance between solar minimum and the combined evolving + solar maximum periods. During solar minimum, and holding other settings fixed, training on NRT-SHARPs generally improves performance for both logistic and LSTM models. In contrast, during the evolving + solar maximum period, NRT-SHARPs decrease performance, particularly in terms of TSS.
The results also demonstrate that models trained on NRT-SHARPs exhibit greater variability than those trained on DEF-SHARPs, likely due to reduced sample quality and a smaller training sample size. In addition, forecasting performance across different forecasting-window lengths is more consistent when models are trained on DEF-SHARPs.
@@ 5 Conclusion
In this paper, we systematically review and sort the most commonly used data products in solar flare forecasting with machine-learning methods. For the flare-response labels, we cross-evaluate the Science-Quality list from NCEI, the operational SWPC-FTP list, and the SSW catalog. We also construct an augmented Science-Quality flare list with assigned NOAA active-region numbers and develop a reproducible processing pipeline for generating this enhanced dataset. For predictor data, we describe the imaging products from AIA/SDO and HMI/SDO and analyze the quality issues that arise near the solar limb. We further examine completeness, missingness, and other quality concerns in the HARP and SHARP parameter datasets.
In addition, we compare the predictive performance of LSTM-based models and logistic-regression baselines across multiple combinations of flare lists (Science-Quality and SWPC-FTP) and SHARP data (DEF-SHARPs and NRT-SHARPs). Our results highlight several important conclusions:
@@
    NRT-SHARPs introduce substantially greater variability in model skill scores, likely resulting from reduced data quality and a smaller effective training sample.
@@
    The impact of data products is strongly solar-cycle dependent. Holding other factors fixed, the operational SWPC-FTP labels improve model performance during solar minimum but degrade performance during solar maximum; NRT-SHARPs exhibit a similar pattern.
@@
By clarifying the strengths and limitations of each product and quantifying their effects on predictive performance, we hope to help researchers in selecting appropriate datasets and to improve the comparability and interpretability of data-driven flare-forecasting studies.
@@ Appendix A Quiet Period Analysis of Active Regions
In this section, we are trying to clearly identify if there are any distributional differences when an active region has an ongoing flare, versus when it is during a “quiet period”. As an initial approach, we are defining the ”quiet period” of an active region to be any period of time that is 24 hours removed from any flare activity (i.e. more than 24 hours before the flare start time and after the flare ends). We are measuring these distributional differences by looking at the marginal distribution of all the SHARPs depending on if it’s a quiet period or a flaring period (which is defined to be any time that is not a quiet period).
After running through all the flares from 2010 to 2024 and categorizing the periods of times of all active regions that existed during this interval as quiet or flaring, we analyzed the marginal distriubtion of the SHARPs during this time. These are visualized here in Figure 18
@@
The quiet (blue) line depicts the SHARP parameters during the quiet period of an AR. The green line depicts the SHARP parameters of active regions that are available in the combined SHARPs data but do not have any flares associated with them in the GOES flare list. For the active regions present in the GOES flare list, we also visualized the distribution of the SHARPs during B-flares to investigate if B-flares could serve as a “‘surrogate” for quiet time. The red curve depicts the exact time interval of a B-flare as described by the GOES flare list, and the purple line expands the interval by 24 hours before and after the start and end times, respectively.
This plot shows that the distribution of a few SHARPs differs depending on whether it is during a quiet period or when there is activity. In particular, the distribution of MEANGAM, MEANSHR, and SHGRT45 suggests that there are indeed differences in the behavior of active regions depending on what time period they are in. To better generalize the behavior of active regions, it may be beneficial to incorporate quiet time data of active regions as well.
Additionally, Figure 19 depicts the class of all flares in the GOES flare list starting from 1996, which encapsulates two solar cycles. During the peak of the solar cycle, we can see that a percentage of B-flares is missing. This indicates that the flare list is not necessarily capturing all B-flares, as they could be obscured by the stronger flares that occur more frequently during the solar cycle. As such, the quiet time imaging data during these solar cycle peaks can also serve as a “representation” of the B-flares during this time.
@@
Additionally, the two dashed lines represent the potential shift that is necessary for the GOES flare labels. As discussed in Section 2, there was an erroneous scaling factor being applied to flares measured prior to the GOES-16 satellite. If we adjust for this scaling, all the flares from before 2019 will be shifted up the length from the “black” line” to the red line. In this scenario, we will have systematically missing low-intensity B-flares in our data, so having the quiet data during this time period can help mitigate any potential negative effects of this.

### v2

@@ 4 Comparison of Data Products from a Flare Prediction Perspective
In this section, we compare the SWPC-FTP operational flare list with the NCEI Science-Quality list by evaluating the numerical skill scores of machine learning models trained on definitive or near-real-time SHARP summary vector parameters. Note that we do not exhaust the available machine learning models and do not aim to optimize reported skill scores here, since, as discussed in the Introduction, (i) when properly trained, commonly used machine learning models do not generate significantly different results, and (ii) the skill scores of machine learning models depend heavily on sample selection and sample split schemes. Therefore, we follow standard practice in selecting the machine learning model and sample split to demonstrate the effects of the various data products on the trained model’s performance.
@@ 4.1 Data Preparation and Model Setting
To assess how data quality affects flare forecasting performance, we consider two sources of flare labels (the SWPC-FTP list and the Science-Quality list) and two versions of predictor data (HMI near-real-time SHARPs and definitive calibrated SHARPs), as shown in Table 9. We evaluate two representative model families: a deep learning architecture (LSTM) and a conventional statistical model (logistic regression with PCA, i.e., principal component analysis). All models are trained on Solar Cycle 24 (2010-05-01 to 2019-12-31) and tested independently on years 2020–2024. The task is formulated as the prediction of $\geq$ M-class flares in the next 24, 12, 6 hours.
@@
We follow the data preparation and sample construction methods described in  (Jiao et al., 2020a). The input data consists of 24-hour time series extracted from the full time series of SHARP parameters of ARs. To ensure data quality, some sequences are excluded, particularly when the ARs are located near the limb. The criteria for excluding unqualified time sequences are as follows.
@@
We focus on a binary classification problem that distinguishes strong(M/X) from Weak (A/B) flares. We exclude the C-class category because its physical significance is ambiguous. C-class flares may be comparable to either small M-class flares or strong B-class flares, making it an unreliable separator boundary and thus vague the training process (Chen et al., 2019b). In this study, we adopt two models: an LSTM model with a sigmoid output layer and a logistic regression model. The LSTM-based model is widely used in solar flare forecasting for handling time-series inputs ((Chen et al., 2019a), (Liu et al., 2019a), (Wang et al., 2020b)) and has been shown to outperform other machine learning models ((Wang et al., 2020b)). The network consists of a two-layer LSTM with a hidden size of 30 units and dropout of 0.3 between layers, followed by a ReLU activation and a fully connected layer that maps the final (or truncated) hidden representation to a single scalar logit. This logit value is interpreted as the pre-sigmoid score for the probability of a strong flare. We include logistic regression due to its simplicity, which in fact results in stability for trained models and interpretability. Given the $k$-hour forecasting window used to extract SHARP parameters, the input of the LSTM model for each target is a $5k\times 20$ matrix, where each row represents a 20-dimensional vector of SHARP parameters. For the logistic model, we first flatten each input along the time dimension into a $100k\times 1$ vector.
@@
    For the logistic regression model, we then apply principal component analysis (PCA) to reduce dimensionality, retaining 98% variance. For the 24-hr forecasting window, 20 principal components already account for over 98% of the variance. The projection coefficients for these principal components are derived from the training set.
@@
    The LSTM model is optimized using the Adam’s optimizer with a binary cross-entropy loss.
@@
    The logistic regression model is fitted with $L_{2}$ regularization. The penalty strength is optimized with 5-fold cross-validation and the ROC-AUC (Receiver Operating Characteristic and Area Under the Curve) scoring objective.
After each bootstrap model is trained, we evaluate probabilistic predictions on the held-out validation set and determine the optimal decision threshold that maximizes the True Skill Statistic (TSS) over a grid of candidate thresholds in $(0,1)$. The optimal threshold and corresponding validation TSS scores are recorded for every bootstrap sample. To evaluate the performance of the machine learning algorithms, we split the samples (time series of SHARP parameters and their corresponding maximum flare magnitudes) into training and test sets. Note that splitting by Active Region (AR) or HARP is crucial to prevent information leakage from the training set to the testing set. In this study, all training/testing splits are performed chronologically by HARP, which mimics the operational setting but may not result in the highest skill scores. See Wang et al. (2020a) for detailed discussions on sample splitting schemes and their impacts on flare prediction results.
It is important to note that varying definitions of positive and negative samples make direct comparisons of results from existing literature challenging, as discussed in Chen et al. (2024). The purpose of this paper is not to propose a universal method for sample construction but rather to compare model performance across different datasets. We also demonstrate numerically that differences arising from data inconsistencies or quality are unaffected by the choice of sample construction method.
@@ 4.2 Science-Quality versus Operational “Response”
To quantify the effect of flare-label quality on the skill scores of the prediction models, we compare an LSTM model and a logistic regression model for 6/12/24 forecasting, trained using Science-Quality and SWPC-FTP flare lists, while using the definitive SHARPs parameter as inputs. Box plots of six commonly adopted skill scores (TSS, HSS, POD, F1, FAR, ACC, see, e.g., Chen et al. (2019a) for definitions) for the 24-hour forecasting windows are shown in Figure 13 and Figure 14 for the LSTM model and the logistic regression model, respectively. The 6-hour and 12-hour window prediction results exhibit consistent trends as these. The numerical results for all model setups are presented in the Tables in Appendix B.
@@
We trained models using data from Solar Cycle 24 (5/1/2010–12/31/2019) and tested them using different time intervals in Solar Cycle 25. The evaluation spans four distinct testing intervals that correspond to different phases of Solar Cycle 25. The 2020–2021 period reflects the early rising phase and is near solar minimum, characterized by extremely low flare occurrence and a highly imbalanced dataset. The 2022 period represents an evolving phase with increasing magnetic complexity and flare frequency. The 2023–2024 interval corresponds to the solar maximum, featuring intense flare activity and a high positive sample ratio (often exceeding 0.8). Finally, the combined 2020–2025 evaluation window averages performance across the full range of solar activity and reflects the model’s overall performance with a balanced test set.
@@
Across the test periods when the solar activity is relatively low, the models trained with SWPC-FTP operational flare list achieve higher predictive skill scores than those trained with Science-Quality labels. This improvement is clearer for logistic models. During the solar minimum (20200101-20211231), the logistic regression model improves from a TSS of 0.15 [0.11,0.18] (Science-Quality) to 0.37 [0.35,0.38] (SWPC-FTP), and the LSTM model improves from 0.42 [0.16,0.59] to 0.57 [0.40,0.69]. Similar margins in performance wins are observed in other metrics. However, for the LSTM model, the difference is not substantial. However, we do observe more stability with the logistic regression model results (with consistently shorter boxes in Figure 14 as compared to Figure 13) as compared to LSTM results. During the high-activity period (2023–2024) when the positive-sample ratio exceeds 0.85, the Science-Quality-trained models achieve better performance except for the TSS score of the logistic models. Across all test periods, models trained with SWPC-FTP lists achieve lower FAR, demonstrating that models trained with operational flare lists are better at identifying flares without increasing false alarms.
@@
Figure 13: Box plots of six skill scores for different testing year choices (given in the x-axis) in the solar cycle 25 of the M/X versus A/B classification task by the LSTM model trained on definitive SHARPs as inputs with a 24-hour forecasting window. The blue boxes represent the Science-Quality flare list, while the orange boxes represent the SWPC-FP operational flare list. The dashed lines indicate the positive sample proportions in the test set, corresponding to the Y axis on the right-hand side of each sub-figure. Note that a smaller FAR means better performance. On each box, the black line marks the mean of the 30 independent bootstrap runs, and the triangle marks the median. The lower and upper bounds of the boxes correspond to the first and third quartiles $Q_{1}$ and $Q_{3}$. The upper and lower error bars are at $Q_{3}+1.5(Q_{3}-Q_{1})$ and $Q_{3}-1.5(Q_{3}-Q_{1})$, respectively. The small dots outside the boxes indicate data points that fall outside the error bars (outliers). The mean value and median are calculated, including the outliers.
@@
During solar minimum, LSTM models trained with the Science-Quality flare lists achieve higher TSS and F1 scores than the logistic regression models trained on the same labels. For the SWPC-FTP lists, the LSTM models perform slightly better or about the same as the logistic models. However, during solar maximum, when trained on identical flare lists, the LSTM models perform worse than the logistic models. In addition, the performance differences between LSTM and logistic models are noticeably larger during solar maximum than during solar minimum. This suggests that the more complex flare environment during solar maximum makes model choice more critical and that simpler models, such as logistic regression, can have an advantage.
@@ 4.3 Near-real-time versus Definitive SHARP
We begin by examining the forecasting model performances when using the DEF-SHARPs and NRT-SHARPs datasets. Although the HARP numbers differ between the two products, we find that the conditional flare-response distributions (given the NOAA AR number) are highly similar. For each SHARP observation (every 12 minutes), we match flares from the Science-Quality list using the NOAA active-region number and flare peak time. Observations associated with M- or X-class flares are labeled as positive, and those associated with A- or B-class flares are labeled as negative, consistent with our model formulation. Figure 16 shows the yearly positive rate and positive-sample counts for both DEF-SHARPs and NRT-SHARPs. Despite fewer valid observations, the NRT-SHARPs product exhibits a positive rate pattern that closely matches that of the definitive product across all training years.
@@
Figure 17 presents the average TSS and F1 score gains across eight model configurations. Because the NRT-SHARPs dataset is extremely imbalanced during 2023–2024, we compare performance between solar minimum and the combined evolving plus solar maximum periods. During solar minimum, and holding other settings fixed, training on NRT-SHARPs generally improves performance for both logistic and LSTM models. In contrast, during the evolving plus solar maximum period, NRT-SHARPs decrease performance, particularly in terms of TSS. The results also demonstrate that models trained on NRT-SHARPs exhibit greater variability than those trained on DEF-SHARPs, likely due to reduced sample quality and a smaller training sample size. In addition, forecasting performance across different forecasting-window lengths is more consistent when models are trained on DEF-SHARPs.
@@ 5 Conclusion
In this paper, we systematically review and compare the most commonly used data products for solar flare forecasting using machine learning methods. For the flare-response labels, we cross-evaluate the Science-Quality list from NCEI, the operational SWPC-FTP list, and the SSW catalog. We also construct an augmented Science-Quality flare list with assigned NOAA active-region numbers and develop a reproducible processing pipeline for generating this enhanced dataset. For predictor data, we describe the imaging products from AIA/SDO and HMI/SDO and analyze the quality issues that arise near the solar limb. We further examine completeness, missingness, and other quality concerns in the HARP and SHARP parameter datasets.
In addition, we compare the predictive performance of LSTM-based models and logistic-regression baselines across multiple combinations of flare lists (Science-Quality and SWPC-FTP) and SHARP data (DEF-SHARPs and NRT-SHARPs). Our analysis results in the following conclusions.
@@
    Near-real-time SHARPs, as compared with definitive SHARPs, introduce substantially greater variability in model skill scores, likely resulting from reduced data quality and a smaller effective training sample.
@@
    The impact of data products is strongly solar-cycle dependent. Holding other factors fixed, the operational SWPC-FTP labels improve model performance during solar minimum but degrade performance during solar maximum. The NRT-SHARPs and DEF-SHARPs comparison exhibits a similar pattern.
@@
By clarifying the strengths and limitations of each product and quantifying their effects on predictive performance, we aim to help researchers select appropriate datasets and to improve the comparability and interpretability of data-driven flare-forecasting studies.
@@ Appendix A Quiet Period Analysis of Active Regions
In this section, we aim to determine whether there are distributional differences between active regions during ongoing flares and during a “quiet period”. As an initial approach, we define the ”quiet period” of an active region as any period 24 hours removed from flare activity (i.e., more than 24 hours before the flare start time and after the flare ends). We measure these distributional differences by examining the marginal distribution of all SHARPs, depending on whether it’s a quiet period or a flaring period (defined as any time that is not a quiet period). After reviewing all flares from 2010 to 2024 and categorizing the periods of all active regions during this interval as quiet or flaring, we analyzed the marginal distribution of SHARPs during this period. These are visualized here in Figure 18
@@
The quiet (blue) line depicts the SHARP parameters during the quiet period of an AR. The green line depicts the SHARP parameters of active regions that are available in the combined SHARPs data but do not have any flares associated with them in the GOES flare list. For the active regions present in the GOES flare list, we also visualized the distribution of the SHARPs during B-flares to investigate if B-flares could serve as a “‘surrogate” for quiet time. The red curve depicts the exact time interval of a B-flare as described by the GOES flare list, and the purple line expands the interval by 24 hours before and after the start and end times, respectively. This plot shows that the distribution of a few SHARPs differs between quiet periods and periods of activity. In particular, the distributions of MEANGAM, MEANSHR, and SHGRT45 suggest that active regions exhibit different behavior across time periods. To better generalize the behavior of active regions, it may be beneficial to incorporate quiet-time data from active regions as well.
Additionally, Figure 19 depicts the class of all flares in the GOES flare list starting from 1996, which encapsulates two solar cycles. At the peak of the solar cycle, a percentage of B-flares is missing. This indicates that the flare list may not capture all B-flares, as they could be obscured by stronger flares that occur more frequently during the solar cycle. As such, quiet-time imaging data from these solar-cycle peaks can also serve as a representation of the B-flares during this period.
@@
Additionally, the two dashed lines represent the potential shift that is necessary for the GOES flare labels. As discussed in Section 2, there was an erroneous scaling factor being applied to flares measured prior to the GOES-16 satellite. If we adjust for this scaling, all the flares from before 2019 will be shifted up the length from the “black” line” to the red line. In this scenario, we will have systematic missingness of low-intensity B-flares in our data; therefore, having quiet data during this period can help mitigate potential negative effects.